OpenStudy (anonymous):

When using the substitution method, what expression for x in the second equation would be substituted in the first equation? 4x – 7y = –9 x + 3y = 12 –3y + 12 3y – 12 –3y – 12 3y + 12

OpenStudy (anonymous):

x+3y=12 x=12-3y = -3y+12 1st one